# Open Biomedical Knowledge Graphs at Scale: Construction, Federation, and AI Agent Access with Samyama Graph Database

### 17 Mar 2026 v2

@@ Abstract
Our contributions are threefold. First, we describe a reproducible ETL pattern for constructing large-scale KGs from heterogeneous public data sources, with cross-source deduplication, batch loading (both Python Cypher and Rust native loaders), and portable snapshot export. Second, we demonstrate cross-KG federation: loading all three snapshots into a single graph tenant enables property-based joins across datasets, answering questions like “For drugs indicated for diabetes, what are their gene targets and which biological pathways do those targets participate in?”—a query that no single KG can answer alone. Third, we introduce schema-driven MCP server generation: each KG automatically exposes typed tools for LLM agents via the Model Context Protocol, enabling natural-language access to graph queries. We evaluate domain-specific MCP tools against text-to-Cypher and standalone GPT-4o on a new BiomedQA benchmark (40 pharmacology questions), achieving 98% accuracy vs. 0% for text-to-Cypher and 75% for standalone GPT-4o.
All data sources are open-license (CC BY 4.0, CC0, OBO, public domain). Snapshots, ETL code, and MCP configurations are publicly available. The combined federated graph (7.9M nodes, 28M edges) loads in approximately 3 minutes from portable snapshots on commodity cloud hardware (AWS g4dn.4xlarge, 62 GB RAM), and cross-KG queries complete in 80 ms–4 s.
@@ 1 Introduction
    Schema-driven AI agent access. Each KG ships with a Model Context Protocol (MCP) (Anthropic, 2024) server configuration with domain-specific tools. We evaluate MCP tools against text-to-Cypher (GPT-4o) and standalone GPT-4o on a new BiomedQA benchmark: MCP tools achieve 98% accuracy (39/40) vs. 0% for text-to-Cypher and 75% for standalone GPT-4o.
@@ 3 Data Sources
Both KGs draw exclusively from open-license public databases. Table 1 summarizes the sources.
@@ 4.1 ETL Pattern
Both KGs follow a common five-phase pattern:
@@ 5.1 Motivation
[…]
@@ 5.3 Federation Mechanism
-- Pathways disrupted by Phase 3 breast cancer drugs
MATCH (ct:ClinicalTrial)-[:STUDIES]->(cond:Condition)
WHERE ct.phase = ’Phase3’ AND cond.name CONTAINS ’Breast’
WITH ct
MATCH (ct)-[:TESTS]->(int:Intervention)
-[:CODED_AS_DRUG]->(drug:Drug)
WITH DISTINCT drug
MATCH (drug)-[:TARGETS]->(prot1:Protein)
MATCH (prot2:Protein)-[:PARTICIPATES_IN]->(pw:Pathway)
WHERE prot1.uniprot_id = prot2.uniprot_id
RETURN pw.name AS pathway,
count(DISTINCT drug.name) AS drugs_targeting
ORDER BY drugs_targeting DESC LIMIT 15
The WHERE prot1.uniprot_id = prot2.uniprot_id clause is the cross-KG bridge—it joins a Protein node from the Clinical Trials snapshot with a Protein node from the Pathways snapshot.
@@ 5.4 Federation Query Patterns
Pattern Clinical Trials $\rightarrow$ $\rightarrow$ Pathways
Drug $\rightarrow$ Pathway Trial $\rightarrow$ Drug (via intervention) Drug target $\rightarrow$ Protein $\rightarrow$ Pathway
Drug $\rightarrow$ GO Process Trial $\rightarrow$ Drug Drug target $\rightarrow$ Protein $\rightarrow$ GOTerm
Drug $\rightarrow$ PPI Network Trial $\rightarrow$ Drug $\rightarrow$ Protein target Target $\rightarrow$ INTERACTS_WITH neighbors
Disease $\rightarrow$ Pathway Gene $\rightarrow$ Disease association Gene $\rightarrow$ Protein $\rightarrow$ Pathway
Adverse Event $\rightarrow$ Pathway Trial $\rightarrow$ AdverseEvent $\rightarrow$ Drug Drug $\rightarrow$ Protein $\rightarrow$ Pathway
@@ 6 Schema-Driven MCP Server Generation
The Pathways KG MCP server exposes 12 domain-specific tools (Table 7).
@@ 7.1 Construction Performance
Snapshot import uses gzip-compressed JSON-lines format. The Pathways KG loads in 3.6 seconds; the Drug Interactions KG in 0.7 seconds; and the Clinical Trials KG, with its 7.8M nodes, loads in 174 seconds. The Drug Interactions KG also has a Rust native loader that constructs the full graph from source files (DrugBank CSV, DGIdb TSV, SIDER TSV) in 928 ms—orders of magnitude faster than the Python HTTP-based ETL. The combined federated graph (7.9M nodes, 28M edges) uses 33 GB RAM (53% of 62 GB available on the AWS instance).
@@ 7.2 Query Performance
Drugs for diabetes $\rightarrow$ gene targets Drug+Path 10 rows 3.6 s
EGFR drugs $\rightarrow$ side effects Drug Int. 10 rows 3.7 s
Cross-KG: PTGS1 drugs + side effects All 3 10 rows 3.7 s
@@ 7.3 BiomedQA Benchmark
We introduce BiomedQA, a benchmark of 40 pharmacology questions across 7 categories over the three federated KGs. We compare three approaches: domain-specific MCP tools (parameterized Cypher templates), text-to-Cypher (GPT-4o generates Cypher from a schema prompt with 3 examples), and standalone GPT-4o (no database access).
@@
GPT-4o standalone 30/40 (75%) 2,474 ms 213
Text-to-Cypher (GPT-4o) 0/40 (0%) 986 ms 548
MCP tools 39/40 (98%) 651 ms 0
The text-to-Cypher result (0%) is striking: GPT-4o generates syntactically plausible Cypher that silently returns empty results due to schema mismatches (hallucinated property names, incorrect multi-MATCH patterns). MCP tools eliminate this class of failure entirely. The benchmark and evaluation scripts are open-source.
@@ 7.4 Federation Correctness
We validated cross-KG federation on the Mac Mini M4 with the following results:
@@
    Bridge construction: 32 breast cancer drug nodes were created from a curated DrugBank mapping, yielding 36 TARGETS edges (Drug $\rightarrow$ Protein) and 199 CODED_AS_DRUG edges (Intervention $\rightarrow$ Drug). The 1,544 matched drug interventions span 5,338 distinct drug names in breast cancer trials.
@@
    Join overlap: 37,990 Protein nodes with uniprot_id from the Pathways KG are available for joining. Of the 36 drug-target proteins, all are present in the Pathways KG’s protein-protein interaction network.
@@
    Federation query result: The 6-hop query “pathways disrupted by breast cancer trial drugs” returns 10 pathways, with Disease (21 drugs), Signal Transduction (19 drugs), Immune System (12 drugs), and Cell Cycle (12 drugs) as the top results. These align with known breast cancer biology: HER2 signaling (Trastuzumab), CDK4/6 cell cycle control (Palbociclib, Ribociclib), and immune checkpoint blockade (Pembrolizumab, Atezolizumab).
@@ 9.2 Limitations
    Memory requirements: The combined 7.8M-node graph requires approximately 8GB RAM. Machines with less memory can load the Pathways KG alone (119K nodes, <<500MB).
@@
    No cross-tenant queries: Currently, federation requires loading both snapshots into a single tenant. Native cross-tenant query support is planned for a future Samyama release.
@@ 10 Conclusion
We presented three open-source biomedical knowledge graphs—Pathways KG (119K nodes from 5 sources), Clinical Trials KG (7.7M nodes from 5 sources), and Drug Interactions KG (33K nodes from 3 sources)—built on Samyama Graph Database. Together they contain 7.9 million nodes and 28 million edges from 13 public data sources. Loading all three snapshots into a single graph tenant enables cross-KG federated queries that bridge molecular biology, translational medicine, and pharmacogenomics. Cross-KG queries complete in 80 ms–4 s on commodity hardware (62 GB RAM). A Rust native loader constructs the Drug Interactions KG in under 1 second.
We introduced domain-specific MCP tools for LLM agent access and evaluated them on a new BiomedQA benchmark (40 pharmacology questions): MCP tools achieve 98% accuracy vs. 0% for text-to-Cypher and 75% for standalone GPT-4o, demonstrating that for domain-specific data access, the LLM’s role should be tool selection and argument extraction, not query generation.

### 18 Mar 2026 v3

@@ Abstract
Our contributions are threefold. First, we describe a reproducible ETL pattern for constructing large-scale KGs from heterogeneous public data sources, with cross-source deduplication, batch loading (both Python Cypher and Rust native loaders), and portable snapshot export. Second, we demonstrate cross-KG federation: loading all three snapshots into a single graph tenant enables property-based joins across datasets, answering questions like “For drugs indicated for diabetes, what are their gene targets and which biological pathways do those targets participate in?”—a query that no single KG can answer alone. Third, we introduce schema-driven MCP server generation: each KG automatically exposes typed tools for LLM agents via the Model Context Protocol, enabling natural-language access to graph queries. We evaluate domain-specific MCP tools against text-to-Cypher and standalone GPT-4o on a new BiomedQA benchmark (40 pharmacology questions), achieving 98% accuracy vs. 85% for schema-aware text-to-Cypher and 75% for standalone GPT-4o, with zero schema errors.
All data sources are open-license (CC BY 4.0, CC0, OBO, public domain). Snapshots, ETL code, and MCP configurations are publicly available. The combined federated graph (7.9M nodes, 28M edges) loads in approximately 3 minutes from portable snapshots on commodity cloud hardware (62 GB RAM), with single-KG queries completing in 80–100 ms and cross-KG federation joins in 1–4 s.
@@ 1 Introduction
    Schema-driven AI agent access. Each KG ships with a Model Context Protocol (MCP) (Anthropic, 2024) server configuration with domain-specific tools. We evaluate MCP tools against text-to-Cypher (GPT-4o) and standalone GPT-4o on a new BiomedQA benchmark: MCP tools achieve 98% accuracy (39/40) vs. 85% for schema-aware text-to-Cypher and 75% for standalone GPT-4o, with zero schema errors.
@@ 3 Data Sources
All three KGs draw exclusively from open-license public databases. Table 1 summarizes the sources.
@@ 4.1 ETL Pattern
All three KGs follow a common five-phase pattern:
@@ 5.1 Motivation
This query requires traversing three KGs: Drug $\xrightarrow{\text{HAS\_INDICATION}}$ Indication (Drug Interactions KG), Drug $\xrightarrow{\text{INTERACTS\_WITH\_GENE}}$ Gene $\xrightarrow{\text{bridge: gene\_name = name}}$ Protein $\xrightarrow{\text{PARTICIPATES\_IN}}$ Pathway (Pathways KG). The WHERE p.name = g.gene_name clause bridges Drug Interactions to Pathways.
@@ 5.3 Federation Mechanism
-- Drug targets -> biological pathways (Drug Interactions -> Pathways)
MATCH (d:Drug {name: ’Metformin’})-[:INTERACTS_WITH_GENE]->(g:Gene)
MATCH (p:Protein)-[:PARTICIPATES_IN]->(pw:Pathway)
WHERE p.name = g.gene_name
RETURN g.gene_name, pw.name LIMIT 10
-- Drug -> clinical trials testing it (Drug Interactions -> Clinical Trials)
MATCH (d:Drug {name: ’Warfarin’})
MATCH (i:Intervention)<-[:TESTS]-(ct:ClinicalTrial)
WHERE i.name = d.name
RETURN ct.nct_id, ct.phase LIMIT 10
-- Breast cancer trial landscape (Clinical Trials only)
MATCH (ct:ClinicalTrial)-[:STUDIES]->(c:Condition)
WHERE c.name CONTAINS ’Breast’
RETURN c.name, count(ct) AS trials
ORDER BY trials DESC LIMIT 5
The WHERE p.name = g.gene_name clause is the Drug Interactions $\to$ Pathways bridge—it joins a Gene node from the Drug Interactions KG with a Protein node from the Pathways KG using gene symbol as the shared identifier. The WHERE i.name = d.name clause bridges Drug Interactions $\to$ Clinical Trials via drug name.
@@ 5.4 Federation Query Patterns
Pattern From KG Bridge To KG
Drug $\rightarrow$ Pathway Drug Int. (Gene) gene_name Pathways (Protein)
Drug $\rightarrow$ Trial Drug Int. (Drug) name Clin. Trials (Intervention)
Drug $\rightarrow$ GO terms Drug Int. (Gene) gene_name Pathways (GOTerm)
Drug SE $\rightarrow$ Pathway Drug Int. (SE+Gene) gene_name Pathways (Protein)
Trial $\rightarrow$ Side effects Clin. Trials (Intervention) name Drug Int. (Drug)
@@ 6 Schema-Driven MCP Server Generation
Each KG’s MCP server exposes domain-specific tools. The Pathways KG has 12 tools (Table 7), the Drug Interactions KG has 12 tools (e.g., drug_interactions, interaction_checker, polypharmacy_risk, drug_side_effects), and the Clinical Trials KG has 15 tools.
@@ 7.1 Construction Performance
Snapshot import uses gzip-compressed JSON-lines format. The Pathways KG loads in 3.4 seconds; the Drug Interactions KG in 0.7 seconds; and the Clinical Trials KG, with its 7.8M nodes, loads in 177 seconds. The Drug Interactions KG also has a Rust native loader that constructs the full graph from source files (DrugBank CSV, DGIdb TSV, SIDER TSV) in 928 ms—orders of magnitude faster than the Python HTTP-based ETL. The combined federated graph (7.9M nodes, 28M edges) uses 33 GB RAM (53% of 62 GB available on the AWS instance).
@@ 7.2 Query Performance
Diabetes drugs $\rightarrow$ pathways Drug Int.+Path 5 rows 3.9 s
Warfarin $\rightarrow$ trial conditions Drug Int.+CT 5 rows 5.8 s
PTGS1 drugs $\rightarrow$ side effects Drug Int. 10 rows 3.7 s
@@ 7.3 BiomedQA Benchmark
We introduce BiomedQA, a benchmark of 40 pharmacology questions across 7 categories over the three federated KGs. We compare three approaches: domain-specific MCP tools (parameterized Cypher templates), text-to-Cypher via the schema-aware NLQ endpoint (GPT-4o with full schema system prompt and few-shot examples), and standalone GPT-4o (no database access). The BiomedQA benchmark is open-source. [^2] https://github.com/samyama-ai/biomedqa
@@
GPT-4o standalone 30/40 (75%) 2,805 ms 195
Text-to-Cypher (NLQ) 34/40 (85%) 1,846 ms 0
MCP tools 39/40 (98%) 920 ms 0
Text-to-Cypher achieves 85% with a schema-aware NLQ pipeline (full schema in system prompt, few-shot examples). Its 6 failures are: 3 schema hallucinations (non-existent edge traversals), 1 exact-vs-CONTAINS mismatch, 1 inline property variable, and 1 correct empty result. MCP tools eliminate schema errors entirely. (Token count is 0 because the NLQ endpoint handles the LLM call server-side.)
@@ 7.4 Federation Correctness
We validated cross-KG federation on the AWS g4dn.4xlarge with all three KGs loaded (7.9M nodes):
@@
    Drug Interactions $\to$ Pathways bridge: The query “Metformin gene targets $\to$ biological pathways” uses WHERE p.name = g.gene_name to bridge Gene nodes from DGIdb to Protein nodes from Reactome/STRING. Returns pathway memberships (e.g., HNF1B $\to$ Developmental Biology) in 3.0 s.
@@
    Drug Interactions $\to$ Clinical Trials bridge: The query “clinical trials testing Warfarin” uses WHERE i.name = d.name to bridge Drug nodes to Intervention nodes. Returns trial NCT IDs and phases (e.g., NCT00835861, PHASE2) in 0.6 s.
@@
    Three-KG chain: The query “drugs indicated for diabetes $\to$ gene targets $\to$ pathways” traverses Drug $\to$ Indication (Drug Int. KG) $\to$ Gene (Drug Int. KG) $\to$ Protein $\to$ Pathway (Pathways KG), returning pathways such as Circadian clock and Dissolution of Fibrin Clot via SERPINE1, in 3.9 s.
@@ 9.2 Limitations
    Memory requirements: The combined 7.9M-node graph requires approximately 33 GB RAM. Machines with less memory can load the Pathways + Drug Interactions KGs alone (151K nodes, <<1 GB).
@@
    No cross-tenant queries: Currently, federation requires loading all snapshots into a single tenant. Native cross-tenant query support is planned for a future Samyama release.
@@ 10 Conclusion
We presented three open-source biomedical knowledge graphs—Pathways KG (119K nodes from 5 sources), Clinical Trials KG (7.7M nodes from 5 sources), and Drug Interactions KG (33K nodes from 3 sources)—built on Samyama Graph Database. Together they contain 7.9 million nodes and 28 million edges from 13 public data sources. Loading all three snapshots into a single graph tenant enables cross-KG federated queries that bridge molecular biology, translational medicine, and pharmacogenomics. Single-KG queries complete in 80–100 ms; cross-KG federation joins in 1–4 s on commodity cloud hardware (62 GB RAM). A Rust native loader constructs the Drug Interactions KG in under 1 second.
We introduced domain-specific MCP tools for LLM agent access and evaluated them on a new BiomedQA benchmark (40 pharmacology questions): MCP tools achieve 98% accuracy vs. 85% for schema-aware text-to-Cypher and 75% for standalone GPT-4o—with zero schema errors, demonstrating that for domain-specific data access, the LLM’s role should be tool selection and argument extraction, not query generation.